# Linear Attention Sequence Parallelism

### v1

@@ 4.1 Scalability and Speed Comparison
TNL  (Qin et al., 2024a) 0 .4B DDP 3.719 LASP + DDP 3.715
Legacy DDP 3.709 LASP + Legacy DDP 3.705
FSDP 3.717 LASP + FSDP 3.714
ZeRO-1 3.653 LASP + ZeRO-1 3.653
ZeRO-2 3.655 LASP + ZeRO-2 3.649
ZeRO-3 3.656 LASP + ZeRO-3 3.649
Legacy DDP 5.425 LASP + Legacy DDP 5.413
FSDP 5.428 LASP + FSDP 5.441
ZeRO-1 5.114 LASP + ZeRO-1 5.118
ZeRO-2 5.105 LASP + ZeRO-2 5.120
ZeRO-3 5.110 LASP + ZeRO-3 5.123
We furthermore conducted a comparison of sequence parallelism on TNL 1B and 7B models against two existing SP methods: DeepSpeed-Ulysses (Jacobs et al., 2023) and Megatron-SP (Korthikanti et al., 2022). All results presented in Fig. 4 are obtained on 64 GPUs. We keep using a fixed batch size of 1, in order to highlight the ability of LASP to handle extremely lone sequences.
LASP demonstrates a notable enhancement in throughput for linear attention, primarily due to its efficient communication design that facilitates the exchange of linear attention intermediate states. Specifically, LASP outperforms DeepSpeed-Ulysses by $38\%$ and Megatron by $136\%$ in terms of throughput at 256K sequence length on 1B model, with the performance gap widening as the sequence length increases. Additionally, system optimizations like kernel fusion and KV State caching enable LASP to support the longest sequence lengths within the same cluster, achieving 2048K for the 1B model and 512K for the 7B model.
4.2 Convergence
Table 2 presents convergence results of two linear-attention based models: TNL and Linear Transformer, evaluated on an epoch-by-epoch basis. The experiments were conducted using the same training corpus: the Pile (Gao et al., 2020). Both linear models has 0.4B parameters, demonstrated consistent loss values when training with or without LASP. All experiments undergoes 50K steps. The uniform loss convergence across various DDP backends demonstrates that LASP does not negatively affect model convergence.
4.3 Ablation on Activation Reducing Methods
LASP prominently reduces the activation memory usage during training process on per GPU, which is orthometric with another activation memory reducing method: activation checkpointing. Following we conduct ablation experiments on AC and LASP to reveal their performance on memory reduction. With pure DDP and FSDP, the maximum sequence lengths are able to train on 8 GPUs are 12K and 16K, respectively. Both AC and LASP can enlarge the maximum sequence length markedly, but encounters slightly throughput reduction. The distinction is the scaling-up performance of LASP is directly proportional to the number of GPUs used. By combining AC and LASP, we can obtain surprising maximum sequence lengths 496K and 768K on single node with using DDP and FSDP backends, respectively.
Table 3: Ablation on Activation Reducing Methods. Both DDP and FSDP backends are tested. A single node equipped with 8 A100 80G GPUs is used to train a TNL-1B model, still with a batch size of 1 for all experiments.
Method Maximum Sequence Length Throughput (tokens/sec)
DDP 12K 131286.0
DDP+AC 64K 117429.5
DDP+LASP 96K 126829.4
DDP+AC+LASP 496K 100837.8
FSDP 16K 145303.6
FSDP+AC 96K 114464.0
FSDP+LASP 120K 138598.8
FSDP+AC+LASP 768K 106578.3
5 Conclusion
We presented LASP, effectively addressing the limitations of existing SP methods on linear transformers by leveraging the specific features of linear attention, which significantly enhanced parallelism efficiency and usability for linear attention-based language models. Through the implementation of an efficient P2P communication mechanism and engineering optimizations such as kernel fusion and KV state caching, LASP achieved a notable reduction in communication traffic and improved hardware utilization on GPU clusters. Compatibility with all types of batch-level DDP methods ensured the practicability of LASP for large-scale distributed training. Our experiments highlighted the advantages of LASP on scalability, speed, memory usage and convergence performance for linear transformers, comparing with existing SP methods in out-of-box frameworks.
Broader Impact
This work presents a significant advancement in the field of artificial intelligence and machine learning, particularly in enhancing the efficiency and scalability of linear attention-based language models. By enabling the processing of sequences up to significantly longer than current methods and accelerating computational speed, LASP has the potential to vastly improve the performance in natural language understanding, genomic sequence analysis, time-series forecasting, and more. However, the increased capability and efficiency may also raise ethical and societal concerns, such as potential misuse in creating persuasive but misleading information, or in surveillance technology. Despite these concerns, the contributions of LASP on reducing computational costs and energy consumption in training large models could have positive environmental implications.
Acknowledgement
This work is partially supported by the National Key R&D Program of China (NO.2022ZD0160100).
Appendix A Appendix
A.1 Hardware and Software
Hardware.
Our experimental configuration involves a maximum of 16 DGX-A100 servers, each equipped with 8 A100 GPUs, these GPUs are interconnected through NVSwitch, ensuring an inter-GPU bandwidth of 600GBps. For inter-node communication, we employ RoCE (RDMA over Converged Ethernet) technology, utilizing 8 RoCE RDMA adapters in each server. This setup facilitates efficient inter-server communication with a bandwidth capacity of 800Gbps.
Software.
Experiments are implemented in PyTorch 2.1.1 and Triton 2.0.0 with CUDA 11.7, cuDNN 8.0, and NCCL 2.14.3. Our algorithm is developed upon Metaseq 0.0.1. The Linear Transformer experiments build on https://github.com/idiap/fast-transformers.
A.2 Additional Experiment Results

### v3

@@ A.4 Generalization of LASP
Cosformer 0.4B DDP 4.001 LASP+DDP 4.005
Cosformer 0.4B ZeRO-1 4.013 LASP+ZeRO-1 3.969
RetNet 0.4B DDP 4.312 LASP+DDP 4.306
RetNet 0.4B ZeRO-1 4.312 LASP+ZeRO-1 4.309
Mamba 0.4B DDP 4.116 LASP+DDP 4.122
Mamba 0.4B ZeRO-1 4.108 LASP+ZeRO-1 4.110
A.5 Additional Experiment Results
A.5.1 Hybrid SP Results on Hybrid Models
We perform a small-scale experiment (8× A100, 1B parameters, DDP backend) to evaluate the feasibility of the hybrid SP approach for hybrid models. In this setup, a “1/4 hybrid” model denotes a configuration where one out of every four layers is a softmax attention Transformer layer. Using “S” to represent softmax attention and “L” for linear attention, a 16-layer “1/4 hybrid” model follows the pattern “LLLSLLLSLLLSLLLS.” The results in Table 7 demonstrate that hybrid SP effectively extends the maximum trainable sequence length for both TNL and Linear Transformer, while incurring only a slight reduction in training speed.
Table 7: Hybrid SP Results on Inter-layer Hybrid Models. “1/4 hybrid” refers to a model where one out of every four layers is a softmax attention Transformer layer. Maximum Sequence Length and Throughput (tokens/sec) are reported.
Model Method Maximum Sequence Length Throughput
1/4 Hybrid TNL No SP 12K 128684
1/4 Hybrid TNL Hybrid SP Solution 90K 125397
1/4 Hybrid Linear Transformer No SP 12K 129253
1/4 Hybrid Linear Transformer Hybrid SP Solution 90K 125883
A.5.2 Evaluation Results on Downstream Tasks
We conduct an experiment with extended training duration of 300K steps (which consumes 40B tokens) to assess the performance of LASP, and its evaluation results on downstream tasks. Both TNL and Linear Transformer with 0.4B parameters are investigated. We evaluate the performance of the trained models on multiple downstream benchmarks, including PIQA, HellaSwag (HS), WinoGrande (WG), ARC-E, ARC-C, OBQA, and CSR-AVG. The results are presented in the Tables 8 and 9. LASP does not negatively affect downstream task performance.
Table 8: Convergence Results of LASP with Extended 300K Steps. Both TNL and Linear Transformer with 0.4B parameters are tested with a batch size of 2 and sequence length of 16K.
Model Parameters Steps Method Loss PPL Method Loss PPL
TNL 0.4B 300K DDP 3.218 9.318 LASP+DDP 3.218 9.321
Linear Transformer 0.4B 300K DDP 4.164 17.972 LASP+DDP 4.145 17.730
Table 9: Evaluation Results on Downstream Tasks. HS: HellaSwag, WG: WinoGrande. A higher score indicates better performance.
Model Method Tokens PIQA HS WG ARC-E ARC-C OBQA CSR-AVG
TNL DDP 40B 55.71 28.21 51.30 28.87 23.72 26.00 35.64
TNL LASP+DDP 40B 54.30 28.17 51.54 31.27 24.06 29.60 36.49
Linear Transformer DDP 40B 52.18 25.68 49.80 26.81 25.60 26.40 34.93
Linear Transformer LASP+DDP 40B 52.18 26.07 49.25 26.22 26.71 27.00 35.44
A.5.3 Quantitative Scalability Results